1. The table below gives the demand schedule of bags of cat food

 

Price

(RM per bag of cat food) | Quantity (bags of cat food per year) | Total Revenue (RM)

5 4

4 8

3 12

2 16

1 20

 

a. Complete the table by calculating the total revenue for each row.


b. Between which prices is demand elastic? Inelastic? Unit elastic?


1
Expert's answer
2022-01-25T08:35:36-0500

1. The table below gives the demand schedule of bags of cat food


 a. Complete the table by calculating the total revenue for each row.

The formula to calculate total revenue is given as:

Revenue = Sales x Average Price of Service or Sales Price

Therefore, the revenue calculation is:

·        R = 5*4 = 20

·        R = 4 * 8 = 32

·        R = 3 * 12 = 36

·        R = 2 * 16 = 32

·        R = 1 * 20 = 20

Thus, the completed table is shown below:


b. Between which prices are demand elastic? Inelastic? Unit elastic?

By assessing the change in total revenue as a result of a change in the price of a product or service, a total revenue test approximates the price elasticity of demand.

Demand is elastic between the price of 3 RM and the price of 5 RM because with an increase in price, the total revenue (demand) decreases from 26 RM to 20 RM. The price is inelastic between the price of 2 RM and 4 RM because, with an increase in price, the revenue (demand) remained constant. There is no point where the unit elastic exists because at no point does the change in price cause a proportional change in the quantity demanded.